OpenStudy (anonymous):

HELP PLZZZZ!!!! @iGreen @BloomLocke367 @rafrasan @sleepyjess @TheSmartOne

OpenStudy (anonymous):

What is the approximate volume of a cylinder with a diameter of 14 meters and a height of 10 meters? Use 3.14 for pi. A. 219.8 m3 B. 439.6 m3 C. 1538.6 m3 D. 6154.4 m3

OpenStudy (anonymous):

The formula is \(\large{\pi\ x\ radius\ x\ height}\)

OpenStudy (sleepyjess):

Do you know the formula for volume of a cylinder?

OpenStudy (anonymous):

lol

OpenStudy (sleepyjess):

AMQ, it is r^2

OpenStudy (anonymous):

Oops, I forgot to put that in...

OpenStudy (anonymous):

(Facepalm)

OpenStudy (anonymous):

3.14 x 12 x 10?

OpenStudy (sleepyjess):

why 12?

OpenStudy (anonymous):

The formula is 3.14 x radius^2 x height

OpenStudy (anonymous):

sorry 14!

OpenStudy (anonymous):

No

OpenStudy (sleepyjess):

why 14?

OpenStudy (anonymous):

That's the diameter...

OpenStudy (anonymous):

The radius is half the diameter, so the radius is...?

OpenStudy (anonymous):

y'all are confusing me:(

OpenStudy (sleepyjess):

I'll let AMQ finish this one with you

OpenStudy (sleepyjess):

or not....

OpenStudy (anonymous):

so 3.14 x

OpenStudy (sleepyjess):

Okay, lets just restart from the beginning :) The formula for volume of a cylinder is \(\pi * r^2 * h\)

OpenStudy (anonymous):

x10

OpenStudy (anonymous):

x 7^2?

OpenStudy (sleepyjess):

Where are you getting *10 from?

OpenStudy (anonymous):

hieght

OpenStudy (sleepyjess):

OH ok, sorry,, blonde moment

OpenStudy (igreen):

\(V = \pi r^2 h\) Plug in 3.14 for pi, 7 for r, and 10 for h.

OpenStudy (sleepyjess):

Ok, so what is the radius?

OpenStudy (igreen):

Then simplify

OpenStudy (sleepyjess):

no lol

OpenStudy (anonymous):

1538.6

OpenStudy (anonymous):

It's C.

OpenStudy (igreen):

*Correct

OpenStudy (anonymous):

C

OpenStudy (anonymous):

??

OpenStudy (sleepyjess):

Yep, C is corect :)

OpenStudy (igreen):

@lslam Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

@iGreen, honey this isn't my main I know. But the poor person acted as if they didn't know how to read. But Jess clearly just said it was C also.

OpenStudy (igreen):

She was confirming her answer.

OpenStudy (sleepyjess):

After the asker had already said it was C

OpenStudy (igreen):

You just shouted out a wrong answer..then shouted out the right answer.

OpenStudy (anonymous):

It doesn't matter anymore, It's C. Move on.